# A Survey of the Self Supervised Learning Mechanisms for Vision Transformers

### v5

@@ 2.4 Comparison of SSL with Transfer Learning
In computer vision, transfer learning (TL) [81] is employed in the development of CNN models to solve a new challenging problem by utilizing the pre-trained model’s weight sharing mechanism especially in the presence of data scarcity [82]. Initially, transfer learning was the dominant approach, but recently, SSL has demonstrated promising results across various applications.
Although TL and SSL are not mutually exclusive techniques, they have a few major differences, as discussed in Table. 3. Moreover, Fig .3 shows the TL and SSL workflow. TL and SL have two main steps: (i) training on a source task and (ii) adapting to a specific target task. The aiming of first step is to optimize the network’s weight parameters to obtain a better starting point.
@@
The challenges associated with TL or SSL especially in complex vision problems are due to several factors, including the volume of pre-training data and class imbalances in the source task. A study by [83] examined this issue across multiple sources and target tasks, yielding the following key insights:
@@
  • SSL offers a more robust performance in transitioning between various types of visual tasks. For example, a model trained using SSL might adapt more easily from facial recognition to gesture recognition. Meanwhile TL may offers better performance when switching between these tasks [84].
@@
Table 3: Summary of the difference between transfer learning and SSL
@@
Pre-training and Fine-tuning Pre-training on a source task first and then fine-tuning on a target task [81] Directly trained on an auxiliary task created from the available data [85]
Task Dependence Presumes that the source and target tasks are correlated [83] Does not assume any particular task; aims for generalized feature learning [86]
Data Requirements Works well with few labeled examples if source task data is similar [87] Needs a huge amount of unlabeled data for feature learning. Afterwards fine-tuned on a relatively smaller labeled dataset [32]
Flexibility Limited to the domain of the source task [83] More flexible; aims to learn features useful for a wide range of tasks [86]
Data Efficiency More data-efficient when a closely related source task is available [87] May require more data for the initial learning phase, but becomes efficient when the learned features are applied to multiple tasks [32]
Computational Overhead Generally lower, especially during the fine-tuning phase [81] Usually higher during the initial training phase to learn from unlabeled data [31]
End Goal The end goal (e.g., specific classification task) is often known beforehand [88] Intension is to learn generalized features that can be applied to various downstream tasks [31]
@@ 3.1 Contrastive Methods
Contrastive Learning (CL) stands as a prominent technique in SSL [89, 31, 32]. Its aim is to capture invariant semantics through pairs of random views, termed Contrastive multi-view coding. CL ensures that global projections of representations are similar for positive samples and dissimilar for negative samples. This section delves into literature reviews on Contrastive Learning applied in transformer frameworks [90].
@@
Li et al. [91] suggested a methodology to enhance domain generalization in mammographic image analysis using contrastive learning. Their approach consists of two stages: first, a multi-style and multi-view contrastive learning pretrainig phase, where CycleGAN and image blending techniques generate diverse mammogram styles to enhance feature robustness; second, a fine-tuning stage for downstream tasks such as mass detection, mass matching, BI-RADS classification, and breast density classification. By learning style-variant representations, their method demonstrated improved performance across multiple mammographic datasets, outperforming existing domain generalization techniques. Wang and co-workers in 2022 [92] introduced the method SRCL that is semantically-relevant contrastive learning for histopathology image analysis. Unlike traditional contrastive learning, SRCL evaluates similarity among instances to identify additional positive pairs by arranging various positive examples that share related visual concepts. This approach enriches the diversity of positive pairs and yields more informative patterns. They employee a hybrid model named CTransPath, in which they combines a CNN with a multi-scale Swin Transformer architecture, their methodology acts as a unified feature extractor, addressing both local and global features during pre-training, seeking to acquire global feature interpretations tailored for activities within the domain of histopathology images. The efficacy of SRCL-pretrained CTransPath was evaluated across five distinct downstream tasks spanning nine publicly available datasets, such as retrieving patches, classifying patches, performing weakly-supervised classification of whole-slide images, detecting mitosis, and segmenting glands in colorectal adenocarcinoma. The findings demonstrated that the visual representations derived from SRCL not just achieved state-of-the-art performance throughout all datasets but it also exhibited increased robustness and transferability compared to both supervised and self-supervised ImageNet pre-training techniques.
@@
To address instability issues during training, Chen and his colleagues in 2021 [93] utilized SSL framework in MOCO, representing an advancement in Contrastive Learning (CL) methodologies. MOCO utilizes dictionaries to gauge similarity between encoded keys and queries. Empirical assessments suggested that self-supervised Transformers within the MOCO framework can demonstrate good performance with relatively fewer inductive biases compared to ImageNet supervised ViT. This approach aims for a better balance between simplicity, accuracy, and scalability, showcasing its superiority over supervised pre-training methods in multiple detection and segmentation tasks. The study proposes replacing the patch projection layer in ViT with fixed random patch projections and suggests removing positional embeddings in SSL setups to address instability.
@@
Mo and his team [94] in 2023 introduced MCVT, which focuses on projecting class tokens to embedding space during the initial or final phases of the ViTs using multi-layer perceptrons. In this methodology, InfoNCE loss is applied to low-level features while ProtoNCE loss is introduced in high-level features, enhancing performance in downstream classification tasks. Extensive experimentation transferring MCVT pre-trained backbones to various downstream tasks confirmed its efficacy across multiple vision-related tasks.
@@
Radford in 2021 [67] proposed a Contrastive Language-Image pre-training (CLIP) framework, which incorporates transformers, learns image representations through natural language supervision. In pre-training, CLIP concurrently trains image and text encoders to accurately predict corresponding pairs within a batch of (image, text) training instances. CLIP discerns among the N * N possible (image, text) pairings within a batch. It achieves this by learning a multi-modal embedding space, optimizing the image and text encoders jointly to increase the cosine similarity of embeddings for the N genuine pairs while decreasing the resemblance of embeddings for the ( N 2̂ - N ) incorrect pairings. Symmetric cross-entropy loss is used in this optimization strategy to adjust these similarity scores. After pre-training, the model uses natural language to reference acquired visual concepts, facilitating zero-shot transfer to downstream tasks, demonstrating substantial transferability to diverse computer vision tasks without the need for task-specific training data. Evaluation encompasses benchmarking across more than 30 diverse computer vision datasets. Remarkably, the model demonstrates substantial transferability to most tasks and frequently matches or surpasses fully supervised baselines without requiring task-specific training data.
@@ 3.1.1 Knowledge Distillation Methods
Literature based Review of Knowledge Distillation Caron et al. [61] in 2021 introduced knowledge distillation into SSL to enhance the training of ViTs. Their method, DINO, comprises two identical networks termed student and teacher networks, each with distinct parameters. Both networks include an encoder or base network, utilizing either ViTs or ResNet50, along with a projection head placed on top of the encoder network. The DINO approach involves generating multiple crops from two perspectives: a local perspective and a global perspective. The global crop perspective is fed to the teacher network, while the local crop perspective is fed to the student network. The local crop is a subset of the global perspective, exhibiting overlap between the two. The teacher network guides updates in the student network, but gradient stop is applied to halt backpropagation in the teacher network. Meanwhile, the teacher network weights are adjusted using the EMA (exponential moving average) of the weights from student network.
@@
Li et al. (2022) [18] proposed Efficient Self-Supervised ViTs (EsViT) to achieve a optimized balance between accuracy and efficiency, reducing the cost in building state-of-the-art SSL vision systems while demonstrating enhanced scaling performance on accuracy versus throughput and model size. They introduced a multi-stage Transformer architecture, combining networks to address the issue of semantic information loss when transitioning from a single-stage transformer to a multi-stage architecture. Employing a distillation learning strategy targeting correspondence between patches, they introduced a new non-contrastive region-matching pre-training task, enhancing the model’s ability to learn intricate regional dependencies and substantially improve the quality of visual representations learned. The Multi-stage ViT segments an input RGB image into individual non-overlapping patches, handling each patch as an individual token in the first stage. Later, the patch merging module concatenates the features of each group and applies a linear layer to reduce the number of tokens. Subsequently, a Transformer with a sparse self-attention module enables interactions among the merged features. This process is repeated multiple times, resulting in a hierarchical representation. EsViT, the proposed model, is validated across various tasks. EsViT achieves 81.3% top-1 accuracy, with the linear evaluation protocol, demonstrating 3.5× parameter efficiency and at least 10× higher throughput compared to previous state-of-the-art models. EsViT leads over its supervised counterpart, Swin Transformers, by outperforming on 17 of 18 datasets in downstream linear classification tasks.
@@
Xie et al. (2021) [95] introduced MoBY, a SSL framework tailored for ViTs. This approach combines MoCo v2 and BYOL approaches, yielding notably high accuracy on the ImageNet-1K dataset. MoBY adopts the key queue, momentum design, and contrastive loss from MoCo v2, and it also take asymmetric data augmentations, asymmetric encoders, and momentum scheduler from BYOL. Employing a knowledge distillation approach, MoBY employees a pair of neural networks (online and target) to enhance mutual learning. Both encoders includes a backbone and a projector head, with the online encoder featuring an additional prediction head, rendering the two encoders non-uniform. Gradients progressively optimizes the online encoder, while the target encoder is a moving average of the online encoder updated through momentum. MoBY is evaluated across diverse downstream tasks, such as object detection and semantic segmentation, showcasing robustness and transferability across varied visual tasks. When utilizing DeiT-S and Swin-Transformers as backbone architectures, MoBY achieves top-1 accuracy rates of 72.8% and 75.0%, respectively, on ImageNet-1K.
@@ 3.1.2 Clustering Methods
Clustering-based methods segment the latent space by applying clustering. This clustering can be performed offline, utilizing the entire dataset, as in PCL [96], DeepCluster v2 [37], and SeLa [76], or online, using mini-batches, as in DINO [61] and SwAV [37]. These approaches enforce consistent clustering assignments of positive pairs using cross-entropy loss. However, previous literature have not taken advantage of this beneficial property in semi-supervised scenarios. Our study aims to fill this gap.
@@
Hemalatha et al. (2025) [97] introduced a novel approach to enhance cervical cancer detection by using self-supervised learning techniques on diverse cell images.
Su et al. (2023) [98] introduced a clustering-based SSL framework tailored for ViTs, aiming to improve the performance of tasks related to object detection and segmentation. Their proposed methodology, Feature-Level SSL (FLSL), works at two levels of semantics: intra-view, which deals with individual image, and inter-view, which considers over an entire dataset. FLSL utilizes a bilevel feature clustering approach, integrating mean-shift clustering intrinsic to transformers for extracting modes as patterns with a k-means-based SSL technique, ensuring semantic coherence of extracted representations both locally and globally. At the first semantic level, FLSL optimizes intra-cluster affinity using a self-attention layer to encourage semantic representations within clusters. On the second tier, semantic representations are cultivated through non-empty k-means clustering, with positive samples identified using a cross-attention layer. Experimental outcomes illustrate the efficacy of FLSL in dense prediction tasks, achieving significant improvements in object detection and instance-level segmentation on MS-COCO, utilizing Mask R-CNN with ViT-S/16 and ViT-S/8 as backbones, respectively.
@@ 3.1.3 Generative Methods
Generative approaches, for example Generative Adversarial Networks (GANs) or Autoencoders, provide a promising avenue for pre-training ViTs. These models not only generate realistic images or reconstruct input images, but also encode meaningful representations of visual content, facilitating enhanced learning of features in subsequent tasks [100].
@@
Heba et al. (2025) [89] suggest that both contrastive learning and in-painting are effective auxiliary tasks for SSL in the medical domain, leading to a 5.26% increase in accuracy for polyp classification compared to other weight initialization methods.
@@
Jie et al. (2025) [101] proposed a self-supervised BM-GAN to learn morphological changes in MRI over the course of a disease. BM-GAN, learns both forward and backward temporal dependencies to interpolate missing time-series data at arbitrary time points. By integrating self-supervised learning and a bidirectional GAN framework, the model ensures realistic reconstruction without labeled data, offering a robust solution for handling irregular and incomplete time-series datasets.
@@
Chen et al. (2020) introduced iGPT, an image-based version of GPT, for self-supervised visual learning. Unlike ViT, which uses patch embedding, iGPT directly resizes and flattens images into lower-resolution sequences, which are then fed into a GPT-2 [102] model for autoregressive pixel prediction. Despite its significant computational cost, iGPT achieves high accuracy in CIFAR-10, surpassing supervised wide ResNet.
@@
Bao et al. (2022) [103] introduced BEIT, that is, bidirectional encoder representation from image transformers is a self-supervised vision representation model. Instead of pixel-wise generation, they adopted MIM as a task to pretrain ViTs in a self-supervised manner. BEIT, grounded on a BERT-style visual transformer, reconstructs masked images in the latent space. The approach involves randomly masking parts of image pieces and predicting the visual tokens relevant to these hidden pieces. The main motivation behind this pre-training technique is to restore the base visual tokens utilizing corrupted image patches. The standard visual Transformer serves as the backbone network. Evaluation using thorough fine-tuning for tasks such as image classification and semantic segmentation demonstrated competitive performance without human annotation.
@@
He et al. (2022) [69] proposed MAE (Masked Autoencoders), in this methodology parts of input images are randomly masked and subsequently reconstructed the pixels that were masked, enabling optimized training of large models that demonstrate strong generalize across downstream tasks. This method utilizes an asymmetric encoder-decoder setup, where the encoder only looks at the visible parts of the image, and a small decoder rebuilds the original image using the learned representation and mask tokens. By hiding a large proportion of the input image, e.g., 75%, MAE creates a challenging and useful self-supervised learning task. This scalable method allows large models to train efficiently and effectively, resulting in powerful models that generalize well. For instance, a standard ViT-Huge model achieved the highest accuracy of (87.8%) among techniques using only ImageNet-1K data. Supervised pre-training was outperformed by transfer learning in downstream task, showing strong potential for scaling behavior.

### v6

@@ 2.4 Comparison of SSL with Transfer Learning
In computer vision, transfer learning (TL) (Yosinski et al. 2014) is employed in the development of CNN models to solve a new challenging problem by utilizing the pre-trained model’s weight sharing mechanism especially in the presence of data scarcity (Hassan et al. 2022). Initially, transfer learning was the dominant approach, but recently, SSL has demonstrated promising results across various applications.
Although TL and SSL are not mutually exclusive techniques, they have some major differences. Fig.3 shows the TL and SSL workflow. TL and SL have two main steps: (i) training on a source task and (ii) adapting to a specific target task. The aiming of first step is to optimize the network’s weight parameters to obtain a better starting point.
@@
The challenges associated with TL or SSL especially in complex vision problems are due to several factors, including the volume of pre-training data and class imbalances in the source task. A study by Yang et al. (2020) examined this issue across multiple sources and target tasks, yielding the following key insights:
@@
  • SSL offers a more robust performance in transitioning between various types of visual tasks. For example, a model trained using SSL might adapt more easily from facial recognition to gesture recognition. Meanwhile TL may offers better performance when switching between these tasks (Ericsson et al. 2021).
@@
Table 3: Comparative summary of transfer learning and self-supervised learning (SSL) based on various aspects including task dependence, data requirements, and computational overhead
@@
Pre-training and Fine-tuning Pre-training on a source task first and then fine-tuning on a target task (Yosinski et al. 2014) Directly trained on an auxiliary task created from the available data (Doersch et al. 2015)
Task Dependence Presumes that the source and target tasks are correlated (Yang et al. 2020) Does not assume any particular task; aims for generalized feature learning (Singh 2018)
Data Requirements Works well with few labeled examples if source task data is similar (Raina et al. 2007) Needs a huge amount of unlabeled data for feature learning. Afterwards fine-tuned on a relatively smaller labeled dataset (Chen et al. 2020)
Flexibility Limited to the domain of the source task (Yang et al. 2020) More flexible; aims to learn features useful for a wide range of tasks (Singh 2018)
Data Efficiency More data-efficient when a closely related source task is available (Raina et al. 2007) May require more data for the initial learning phase, but becomes efficient when the learned features are applied to multiple tasks (Chen et al. 2020)
Computational Overhead Generally lower, especially during the fine-tuning phase (Yosinski et al. 2014) Usually higher during the initial training phase to learn from unlabeled data (He et al. 2020)
End Goal The end goal (e.g., specific classification task) is often known beforehand (Oquab et al. 2014) Intension is to learn generalized features that can be applied to various downstream tasks (He et al. 2020)
@@ 3.1 Contrastive Methods
Contrastive Learning (CL) stands as a prominent technique in SSL (El-Shimy et al. 2025, He et al. 2020, Chen et al. 2020). Its aim is to capture invariant semantics through pairs of random views, termed Contrastive multi-view coding. CL ensures that global projections of representations are similar for positive samples and dissimilar for negative samples. This section delves into literature reviews on Contrastive Learning applied in transformer frameworks (Jia et al. 2025).
@@
Li et al. (2025) suggested a methodology to enhance domain generalization in mammographic image analysis using contrastive learning. Their approach consists of two stages: first, a multi-style and multi-view contrastive learning pretrainig phase, where CycleGAN and image blending techniques generate diverse mammogram styles to enhance feature robustness; second, a fine-tuning stage for downstream tasks such as mass detection, mass matching, BI-RADS classification, and breast density classification. By learning style-variant representations, their method demonstrated improved performance across multiple mammographic datasets, outperforming existing domain generalization techniques. Wang et al. (2022) introduced the method SRCL that is semantically-relevant contrastive learning for histopathology image analysis. Unlike traditional contrastive learning, SRCL evaluates similarity among instances to identify additional positive pairs by arranging various positive examples that share related visual concepts. This approach enriches the diversity of positive pairs and yields more informative patterns. They employee a hybrid model named CTransPath, in which they combines a CNN with a multi-scale Swin Transformer architecture, their methodology acts as a unified feature extractor, addressing both local and global features during pre-training, seeking to acquire global feature interpretations tailored for activities within the domain of histopathology images. The efficacy of SRCL-pretrained CTransPath was evaluated across five distinct downstream tasks spanning nine publicly available datasets, such as retrieving patches, classifying patches, performing weakly-supervised classification of whole-slide images, detecting mitosis, and segmenting glands in colorectal adenocarcinoma. The findings demonstrated that the visual representations derived from SRCL not just achieved state-of-the-art performance throughout all datasets but it also exhibited increased robustness and transferability compared to both supervised and self-supervised ImageNet pre-training techniques.
@@
To address instability issues during training, Chen* et al. (2021) utilized SSL framework in MOCO, representing an advancement in Contrastive Learning (CL) methodologies. MOCO utilizes dictionaries to gauge similarity between encoded keys and queries. Empirical assessments suggested that self-supervised Transformers within the MOCO framework can demonstrate good performance with relatively fewer inductive biases compared to ImageNet supervised ViT. This approach aims for a better balance between simplicity, accuracy, and scalability, showcasing its superiority over supervised pre-training methods in multiple detection and segmentation tasks. The study proposes replacing the patch projection layer in ViT with fixed random patch projections and suggests removing positional embeddings in SSL setups to address instability.
@@
Mo et al. (2023) in 2023 introduced MCVT, which focuses on projecting class tokens to embedding space during the initial or final phases of the ViTs using multi-layer perceptrons. In this methodology, InfoNCE loss is applied to low-level features while ProtoNCE loss is introduced in high-level features, enhancing performance in downstream classification tasks. Extensive experimentation transferring MCVT pre-trained backbones to various downstream tasks confirmed its efficacy across multiple vision-related tasks.
@@
Radford et al. (2021) proposed a Contrastive Language-Image pre-training (CLIP) framework, which incorporates transformers, learns image representations through natural language supervision. In pre-training, CLIP concurrently trains image and text encoders to accurately predict corresponding pairs within a batch of (image, text) training instances. CLIP discerns among the N * N possible (image, text) pairings within a batch. It achieves this by learning a multi-modal embedding space, optimizing the image and text encoders jointly to increase the cosine similarity of embeddings for the N genuine pairs while decreasing the resemblance of embeddings for the ( N 2̂ - N ) incorrect pairings. Symmetric cross-entropy loss is used in this optimization strategy to adjust these similarity scores. After pre-training, the model uses natural language to reference acquired visual concepts, facilitating zero-shot transfer to downstream tasks, demonstrating substantial transferability to diverse computer vision tasks without the need for task-specific training data. Evaluation encompasses benchmarking across more than 30 diverse computer vision datasets. Remarkably, the model demonstrates substantial transferability to most tasks and frequently matches or surpasses fully supervised baselines without requiring task-specific training data.
@@ 3.1.1 Knowledge Distillation Methods
Literature based Review of Knowledge Distillation Caron et al. (2021) in 2021 introduced knowledge distillation into SSL to enhance the training of ViTs. Their method, DINO, comprises two identical networks termed student and teacher networks, each with distinct parameters. Both networks include an encoder or base network, utilizing either ViTs or ResNet50, along with a projection head placed on top of the encoder network. The DINO approach involves generating multiple crops from two perspectives: a local perspective and a global perspective. The global crop perspective is fed to the teacher network, while the local crop perspective is fed to the student network. The local crop is a subset of the global perspective, exhibiting overlap between the two. The teacher network guides updates in the student network, but gradient stop is applied to halt backpropagation in the teacher network. Meanwhile, the teacher network weights are adjusted using the EMA (exponential moving average) of the weights from student network.
@@
citeli2022exploring proposed Efficient Self-Supervised ViTs (EsViT) to achieve a optimized balance between accuracy and efficiency, reducing the cost in building state-of-the-art SSL vision systems while demonstrating enhanced scaling performance on accuracy versus throughput and model size. They introduced a multi-stage Transformer architecture, combining networks to address the issue of semantic information loss when transitioning from a single-stage transformer to a multi-stage architecture. Employing a distillation learning strategy targeting correspondence between patches, they introduced a new non-contrastive region-matching pre-training task, enhancing the model’s ability to learn intricate regional dependencies and substantially improve the quality of visual representations learned. The Multi-stage ViT segments an input RGB image into individual non-overlapping patches, handling each patch as an individual token in the first stage. Later, the patch merging module concatenates the features of each group and applies a linear layer to reduce the number of tokens. Subsequently, a Transformer with a sparse self-attention module enables interactions among the merged features. This process is repeated multiple times, resulting in a hierarchical representation. EsViT, the proposed model, is validated across various tasks. EsViT achieves 81.3% top-1 accuracy, with the linear evaluation protocol, demonstrating 3.5× parameter efficiency and at least 10× higher throughput compared to previous state-of-the-art models. EsViT leads over its supervised counterpart, Swin Transformers, by outperforming on 17 of 18 datasets in downstream linear classification tasks.
@@
Xie et al. (2021) introduced MoBY, a SSL framework tailored for ViTs. This approach combines MoCo v2 and BYOL approaches, yielding notably high accuracy on the ImageNet-1K dataset. MoBY adopts the key queue, momentum design, and contrastive loss from MoCo v2, and it also take asymmetric data augmentations, asymmetric encoders, and momentum scheduler from BYOL. Employing a knowledge distillation approach, MoBY employees a pair of neural networks (online and target) to enhance mutual learning. Both encoders includes a backbone and a projector head, with the online encoder featuring an additional prediction head, rendering the two encoders non-uniform. Gradients progressively optimizes the online encoder, while the target encoder is a moving average of the online encoder updated through momentum. MoBY is evaluated across diverse downstream tasks, such as object detection and semantic segmentation, showcasing robustness and transferability across varied visual tasks. When utilizing DeiT-S and Swin-Transformers as backbone architectures, MoBY achieves top-1 accuracy rates of 72.8% and 75.0%, respectively, on ImageNet-1K.
@@ 3.1.2 Clustering Methods
Clustering-based methods segment the latent space by applying clustering. This clustering can be performed offline, utilizing the entire dataset, as in PCL (Li et al. 2020), DeepCluster v2 (Caron et al. 2020), and SeLa (Asano et al. 2019), or online, using mini-batches, as in DINO( (Caron et al. 2021) and SwAV ((Caron et al. 2020). These approaches enforce consistent clustering assignments of positive pairs using cross-entropy loss. However, previous literature have not taken advantage of this beneficial property in semi-supervised scenarios. Our study aims to fill this gap.
@@
Hemalatha and Vetriselvi (2025) introduced a novel approach to enhance cervical cancer detection by using self-supervised learning techniques on diverse cell images.
Su et al. (2024) introduced a clustering-based SSL framework tailored for ViTs, aiming to improve the performance of tasks related to object detection and segmentation. Their proposed methodology, Feature-Level SSL (FLSL), works at two levels of semantics: intra-view, which deals with individual image, and inter-view, which considers over an entire dataset. FLSL utilizes a bilevel feature clustering approach, integrating mean-shift clustering intrinsic to transformers for extracting modes as patterns with a k-means-based SSL technique, ensuring semantic coherence of extracted representations both locally and globally. At the first semantic level, FLSL optimizes intra-cluster affinity using a self-attention layer to encourage semantic representations within clusters. On the second tier, semantic representations are cultivated through non-empty k-means clustering, with positive samples identified using a cross-attention layer. Experimental outcomes illustrate the efficacy of FLSL in dense prediction tasks, achieving significant improvements in object detection and instance-level segmentation on MS-COCO, utilizing Mask R-CNN with ViT-S/16 and ViT-S/8 as backbones, respectively.
@@ 3.1.3 Generative Methods
Generative approaches, for example Generative Adversarial Networks (GANs) or Autoencoders, provide a promising avenue for pre-training ViTs. These models not only generate realistic images or reconstruct input images, but also encode meaningful representations of visual content, facilitating enhanced learning of features in subsequent tasks in (Showrov et al. 2024).
@@
El-Shimy et al. (2025) suggest that both contrastive learning and in-painting are effective auxiliary tasks for SSL in the medical domain, leading to a 5.26% increase in accuracy for polyp classification compared to other weight initialization methods.
@@
Lin et al. (2025) proposed a self-supervised BM-GAN to learn morphological changes in MRI over the course of a disease. BM-GAN, learns both forward and backward temporal dependencies to interpolate missing time-series data at arbitrary time points. By integrating self-supervised learning and a bidirectional GAN framework, the model ensures realistic reconstruction without labeled data, offering a robust solution for handling irregular and incomplete time-series datasets.
@@
Chen et al. (2020) introduced iGPT, an image-based version of GPT, for self-supervised visual learning. Unlike ViT, which uses patch embedding, iGPT directly resizes and flattens images into lower-resolution sequences, which are then fed into a GPT-2 (Radford et al. 2019) model for autoregressive pixel prediction. Despite its significant computational cost, iGPT achieves high accuracy in CIFAR-10, surpassing supervised wide ResNet.
@@
Bao et al. (2021) introduced BEIT, that is, bidirectional encoder representation from image transformers is a self-supervised vision representation model. Instead of pixel-wise generation, they adopted MIM as a task to pretrain ViTs in a self-supervised manner. BEIT, grounded on a BERT-style visual transformer, reconstructs masked images in the latent space. The approach involves randomly masking parts of image pieces and predicting the visual tokens relevant to these hidden pieces. The main motivation behind this pre-training technique is to restore the base visual tokens utilizing corrupted image patches. The standard visual Transformer serves as the backbone network. Evaluation using thorough fine-tuning for tasks such as image classification and semantic segmentation demonstrated competitive performance without human annotation.
@@
He et al. (2022) proposed MAE (Masked Autoencoders), in this methodology parts of input images are randomly masked and subsequently reconstructed the pixels that were masked, enabling optimized training of large models that demonstrate strong generalize across downstream tasks. This method utilizes an asymmetric encoder-decoder setup, where the encoder only looks at the visible parts of the image, and a small decoder rebuilds the original image using the learned representation and mask tokens. By hiding a large proportion of the input image, e.g., 75%, MAE creates a challenging and useful self-supervised learning task. This scalable method allows large models to train efficiently and effectively, resulting in powerful models that generalize well. For instance, a standard ViT-Huge model achieved the highest accuracy of (87.8%) among techniques using only ImageNet-1K data. Supervised pre-training was outperformed by transfer learning in downstream task, showing strong potential for scaling behavior.